# Mechanical Amorphization of Glass-Forming Systems Induced by Oscillatory Deformation: The Energy Absorption and Efficiency Control

### v2

@@ II.1 The initial sample preparation
Three types of glass-forming alloy systems: CuZr, CuZr2, Al3Sm are used to investigate the transition of the crystal to the amorphous state induced by oscillatory deformation. The atomic interactions of all the systems are represented by the potential of the embedded atom method (EAM), for the Cu-Zr systemMendelev2019development and the Al-Sm systemSong2021molecular , respectively. The intermetallic single crystal structure is generated by the Atomsk packageHirel2015Atomsk . For the CuZr system, the single crystal structure consists of 54,000 atoms with the B2 phase, and for the CuZr2 system, the single crystal structure consists of 54,000 atoms with $C11_{b}$ phase, and for Al3Sm system, the single crystal structure consists of 56,320 atoms with the phase $D0_{19}$. All systems were equilibrated in the constant pressure-constant temperature (NPT) ensemble at 300 K for 100 ps.
@@ II.3 The loading process of oscillatory deformation
In this work, oscillatory deformations along three directions are used for mechanical amorphization. The main direction is loading a sinusoidal strain as $\gamma_{A}sin(t/t_{p})$, while the other directions used an opposite direction sinusoidal strain with half the strain amplitude, as $-\gamma_{A}/2sin(t/t_{p})$. The primary coordinates axes and loading directions (tensile and compressive) can both be altered. Hence, for one configuration, there will be six deformation types for statistics. For small strain amplitude deformation, the loading process is equivalent to pure shear deformation with conversation volume, and for large strain amplitude, the deformation can involve a density change, but the average process of volume change is not significant, and this kind of deformation is widely used in simulation to mimic mechanical alloyingRogachev2022MechanicalAI ; Li2024stress and ultrasonic loadingshang2023influence ; Li2023ultrafast . The temperature in all simulations is maintained at 300 K by the Nosé-Hoover thermostatnose1984unified .
@@ III.2 The crystal to amorphous transition (CTA)
Figure 3: Control Parameters of Oscillatory Deformation a, b Evolution of energy absorption ($\Delta U$) and crystal content ($f_{c}$) over loading time for varying loading amplitudes ($\gamma_{A}$) and periods ($t_{p}$), respectively, in the CuZr2 system. The mechanical amorphization time ($t_{a}$) is indicated by arrows, and a dashed line is provided as a guide for the eye. c, d Evolution of energy absorption ($\Delta U_{s}$) and crystal content ($f_{c}$) over loading time for different material systems-CuZr, CuZr2, Al3Sm-under the same loading conditions: strain amplitude $\gamma_{A}=0.2$ and loading period $t_{p}=100$ ps.
@@ III.3 The effect of loading conditions and materials property on CTA
The steady-state energy absorption $\Delta U_{s}$ increases with increasing strain amplitude ($\gamma_{A}$) and decreases with increasing loading period ($t_{p}$) (Figure 4d). In contrast to $t_{a}$, $\Delta U_{s}$ exhibits a similar trend with respect to the strain amplitude between different material systems, without any crossover behavior. As the strain amplitude increases, $\Delta U_{s}$ tends to increase, indicating that larger strain amplitudes generally require more energy to induce amorphization. However, the magnitude of $\Delta U_{s}$ can vary significantly between materials, reflecting differences in their crystallographic structures and microstructural properties. This suggests that while the absorbed energy $\Delta U_{s}$ tends to be higher for larger strain amplitudes, the final energy state can differ significantly between materials, potentially affecting their overall mechanical properties and behavior.
@@ III.4 The control parameters of CTA: work rate and strain rate
Where $t_{0}\equiv 1$ ps for dimensional balance, $\Delta G$ is the free energy barrier for the crystal-to-amorphous transition in the absence of loading, and $T_{\text{e}}\equiv T+\eta t_{0}\left(\frac{\gamma_{A}^{2}}{t_{p}}\right)$ is the effective temperature. The effective temperature consists of two contributions: one from the system’s temperature and the other induced by the external loading, which is assumed to be linearly dependent on the work rate $\gamma_{A}^{2}/t_{p}$. Here $\eta$ is a material-dependent parameter that characterizes the rate of change in the effective temperature with the increasing work rate.
@@
Moreover, $\eta$ shows a linear relationship with the quiescent amorphous energy $\Delta U_{0}$ and a positive correlation with the energy barrier $\Delta G$ (Figures 6(d) and A3). This suggests that the mechanical temperature unit is intricately determined by the properties of the material, independent of the energy state.
At low work rates, the efficiency of mechanical amorphization is mainly controlled by the energy barrier $\Delta G$. Since $\Delta G$ is approximately proportional to the melting enthalpy $\Delta H_{m}$, glass formation systems with smaller $\Delta H_{m}$ exhibit shorter mechanical amorphization times. At higher work rates, the effective temperature increases significantly, dominating the efficiency of the process. In this regime, $\eta\sim\Delta U_{0}$, and materials with larger quiescent amorphous energies $\Delta U_{0}$ tend to have shorter mechanical amorphization times. Interestingly, there is also a positive correlation between $\Delta H_{m}$ and $\Delta U_{0}$ (as well as between $\Delta G$ and $\eta$, see Figures A3 and A4).
@@ V Acknowledgments
The authors thank Prof. Haiyang Bai, Prof. Jean-Louis Barrat, and Prof. Mo Li for insightful discussion. This work is supported by the Pearl River Talent Recruitment Program (Grant No.2021QN02C04), the NSF of China (Grant Nos.12474187, 52130108, 52201176) and Young Talent Support Project of Guangzhou Association for Science and Technology (Grant No.QT2024–041).
@@
Figure 1: Graphical Abstract: (a) the schematic of amorphization efficiency and absorption energy on the two-state potential landscape. (b) amorphization efficiency controlled by work rate. (c) absorption energy controlled by strain rate.

### v3

@@ Abstract
##footnotetext: These authors contributed equally to this work
@@ II.1 The initial sample preparation
Three types of glass-forming alloy systems: CuZr, CuZr2, Al3Sm are used to investigate the transition of the crystal to the amorphous state induced by oscillatory deformation. The Al-Sm and Cu-Zr systems were selected for two principal reasons: (i) to ensure generalizability by investigating the amorphization mechanism across materials with distinct crystal symmetries, glass-forming abilities, and intrinsic properties; and (ii) to address the scarcity of suitable interatomic potentials, as these systems are well-characterized and exhibit good glass-forming ability. The atomic interactions of all the systems are represented by the potential of the embedded atom method (EAM), for the Cu-Zr systemMendelev2019development and the Al-Sm systemSong2021molecular , respectively. The intermetallic single crystal structure is generated by the Atomsk packageHirel2015Atomsk . For the CuZr system, the single crystal structure consists of 54,000 atoms with the B2 phase, and for the CuZr2 system, the single crystal structure consists of 54,000 atoms with $C11_{b}$ phase, and for Al3Sm system, the single crystal structure consists of 56,320 atoms with the phase $D0_{19}$. Those crystal structure and pair distribution functions (PDFs) are shown in FigureA1. All systems were equilibrated in the constant pressure-constant temperature (NPT) ensemble at 300 K for 100 ps.
@@ II.3 The loading process of oscillatory deformation
In this work, oscillatory deformations along three directions are used for mechanical amorphization. The main direction is loading a sinusoidal strain as $\gamma_{A}sin(t/t_{p})$, while the other directions used an opposite direction sinusoidal strain with half the strain amplitude, as $\gamma_{A}/2sin(t/t_{p})$. The primary coordinates axes and loading directions (tensile and compressive) can both be altered. Hence, for one configuration, there will be six deformation types for statistics. For small strain amplitude deformation, the loading process is equivalent to pure shear deformation with conversation volume, and for large strain amplitude, the deformation can involve a density change, but the average process of volume change is not significant, and this kind of deformation is widely used in simulation to mimic mechanical alloyingRogachev2022MechanicalAI ; Li2024stress and ultrasonic loadingshang2023influence ; Li2023ultrafast . The temperature in all simulations is maintained at 300 K by the Nosé-Hoover thermostatnose1984unified .
II.4 The melting enthalpy $\Delta H_{m}$
The melting enthalpy, $\Delta H_{m}$, which represents the enthalpy difference between the liquid and crystalline states at the melting temperature, was calculated for each material as follows. Single-crystal samples were heated from 300 K to 2500 K under an isothermal-isobaric (NPT) ensemble, employing a Nosé–Hoover thermostatnose1984unified to control temperature. A constant heating rate of 0.1 K/ps was applied throughout the simulation. The transition from crystal to liquid was characterized by a discontinuous jump in enthalpy; the magnitude of this jump corresponds to $\Delta H_{m}$. For comparison, additional simulations were performed by quenching the system from 2500 K to 300 K at the same cooling rate of 0.1 K/ps to obtain the corresponding cooling curves.
@@ III.2 The crystal to amorphous transition (CTA)
Figure 3: Control Parameters of Oscillatory Deformation a, b Evolution of energy absorption ($\Delta U$) and crystal content ($f_{c}$) over loading time for varying loading amplitudes ($\gamma_{A}$) and periods ($t_{p}$), respectively, in the CuZr2 system. Note the strain rate $\gamma/t_{p}$ is fixed at 0.001/ps. The mechanical amorphization time ($t_{a}$) is indicated by arrows, and a dashed line is provided as a guide for the eye. c, d Evolution of energy absorption ($\Delta U_{s}$) and crystal content ($f_{c}$) over loading time for different material systems-CuZr, CuZr2, Al3Sm-under the same loading conditions: strain amplitude $\gamma_{A}=0.2$ and loading period $t_{p}=100$ ps.
@@ III.3 The effect of loading conditions and materials property on CTA
The steady-state energy absorption $\Delta U_{s}$ increases with increasing strain amplitude ($\gamma_{A}$) and decreases with increasing loading period ($t_{p}$) (Figure 4c,d). In contrast to $t_{a}$, $\Delta U_{s}$ exhibits a similar trend with respect to the strain amplitude between different material systems, without any crossover behavior. As the strain amplitude increases, $\Delta U_{s}$ tends to increase, indicating that larger strain amplitudes generally require more energy to induce amorphization. However, the magnitude of $\Delta U_{s}$ can vary significantly between materials, reflecting differences in their crystallographic structures and microstructural properties. This suggests that while the absorbed energy $\Delta U_{s}$ tends to be higher for larger strain amplitudes, the final energy state can differ significantly between materials, potentially affecting their overall mechanical properties and behavior.
@@ III.4 The control parameters of CTA: work rate and strain rate
Where $t_{0}\equiv 1$ ps for dimensional balance, $\Delta G$ is the free energy barrier for the crystal-to-amorphous transition in the absence of loading, and $T_{\text{e}}\equiv T+\eta t_{0}\left(\frac{\gamma_{A}^{2}}{t_{p}}\right)$ is the effective temperature. The effective temperature consists of two contributions: one from the system’s temperature and the other induced by the external loading, which is assumed to be linearly dependent on the work rate $\gamma_{A}^{2}/t_{p}$. Note this linear dependence between the effective temperature and the work rate is merely an approximation derived from a Taylor expansion. Here $\eta$ is a material-dependent parameter that characterizes the rate of change in the effective temperature with the increasing work rate.
@@
The measured mechanical response is well-captured by a Herschel-Bulkley constitutive relation. However, while this model provides an excellent empirical fit, the physical origin of the observed strain-rate scaling exponent of n = 1/2 warrants further discussion. This specific value may not be arbitrary but could emerge from the underlying micromechanics of the amorphous phase, particularly the dynamics of Shear Transformation Zones (STZs)PhysRevE.57.7192 . We hypothesize that the exponent of 1/2 is linked to a fundamental scaling between the applied strain rate and the activation energy for STZ operation. In amorphous solids, plastic flow is often governed by a stress-dependent energy barrier, $\Delta E\sim 1/\tau$, where $\tau$ is the shear stressFalk2011deformation . If the effective stress governing STZ activation scales with the square root of the strain rate ($\tau_{\text{eff}}\sim\dot{\gamma}^{1/2}$), it would naturally yield a Herschel-Bulkley exponent of n = 1/2. This scaling could arise from a mean-field interaction between STZs or from a specific form of cooperative, diffusive-like motion within the shear bands. This proposed connection remains speculative and highlights a critical area for future investigation.
Furthermore, our results demonstrate that the steady-state energy absorption, $\Delta U_{s}$, varies significantly across different material systems. This variation is primarily attributed to the intrinsic material property $\Delta U_{s}$, the energy offset between crystalline and amorphous states, which correlates with the melting enthalpy, $\Delta H_{m}$ (see Fig.A5). This suggests that the fundamental thermodynamic stability of the crystal relative to the glass plays a key role in determining how much energy a material can absorb during mechanical amorphization.
However, a direct correlation between traditional microstructural order parameters, such as bond orientation order, and $\Delta U_{s}$ remains elusive. For instance, the crystal fraction ($f_{c}$) exhibits trends divergent from $\Delta U_{s}$ and is itself invariant to loading parameters in the amorphous state (Figure 3c,d). The precise microstructural features that govern energy absorption thus present a compelling open question for future research, potentially requiring more sophisticated descriptors of amorphous structure.
@@
Moreover, $\eta$ shows a linear relationship with the quiescent amorphous energy $\Delta U_{0}$ and a positive correlation with the energy barrier $\Delta G$ (Figures 6(d) and A4). This suggests that the mechanical temperature unit is intricately determined by the properties of the material, independent of the energy state.
At low work rates, the efficiency of mechanical amorphization is mainly controlled by the energy barrier $\Delta G$. Since $\Delta G$ is approximately proportional to the melting enthalpy $\Delta H_{m}$, glass formation systems with smaller $\Delta H_{m}$ exhibit shorter mechanical amorphization times. At higher work rates, the effective temperature increases significantly, dominating the efficiency of the process. In this regime, $\eta\sim\Delta U_{0}$, and materials with larger quiescent amorphous energies $\Delta U_{0}$ tend to have shorter mechanical amorphization times. Interestingly, there is also a positive correlation between $\Delta H_{m}$ and $\Delta U_{0}$ (as well as between $\Delta G$ and $\eta$, see Figures A4 and A5).
@@
Figure 7: A schematic of free energy landscape for amorphous solid, liquid and crystal. In the panel, a disorder phase (liquid or amorphous) grows in the crystal phase.
The relationship between melting enthalpy, $\Delta H_{m}$, and the mechanical amorphization energy barrier, $\Delta G$, is illustrated in Figure 7. In good glass-forming materials, both mechanical amorphization and thermal melting can be viewed as processes involving the nucleation of a disordered phase within an ordered crystalline matrix. The observed positive correlation between $\Delta G$ and $\Delta H_{m}$ can be rationalized within the framework of classical nucleation theory (CNT), wherein the free energy barrier is governed by two competing contributions: the volumetric thermodynamic driving force (which promotes the phase transformation and is related to $\Delta H_{m}$) and the interfacial energy penalty (which opposes it).
The interfacial energy term plays a dominant role in this correlation. In a highly stable crystal—characterized by a large $\Delta H_{m}$, forming a disordered interface requires breaking strong, ordered atomic bonds, resulting in a high energy cost. Thus, materials with greater $\Delta H_{m}$ exhibit higher interfacial energies. The positive dependence of $\Delta G$ on $\Delta H_{m}$ (see Figure 6) indicates that the increased interfacial energy penalty in such stable crystals outweighs the effect of the larger thermodynamic driving force. This highlights that kinetic stability against amorphization is significantly influenced by the difficulty of interface formation.
@@ V Acknowledgments
The authors thank Prof. Haiyang Bai, Prof. Jean-Louis Barrat, and Prof. Mo Li for insightful discussion.
VI Funding
This work is supported by the Pearl River Talent Recruitment Program (Grant No.2021QN02C04), the NSF of China (Grant Nos.12474187, 52130108, 52201176) and Young Talent Support Project of Guangzhou Association for Science and Technology (Grant No.QT2024–041).
VII Author contributions
B.S.S. and X.X.L contributed equally to this work. B.S.S, P.F.G and W.H.W. conceived and supervised this work; B.S.S designed and conducted all the simulations with assistance from X.X.L; B.S.S., X.X.L., and P.F.G. wrote the paper with input and comments from all authors.
@@ VII Author contributions
Figure A1: Crystal structures of the three model systems: a CuZr, b CuZr2, and c Al3Sm. The corresponding radial distribution functions (RDFs) calculated at 300 K are presented in panels d, e, and f, respectively.